OpenStudy (anonymous):

Write the equation of the line that passes through the point (6, -2) and has a slope of -2/3.

OpenStudy (turingtest):

a line with slope \(m\) that passes through the point \((x_0,y_0)\) has the equation \[y-y_0=m(x-x_0)\] what is the slope in your problem? what are \(x_0\) and \(y_0\) ?

OpenStudy (anonymous):

y-(-2) = -2/3(x-6) ?

OpenStudy (anonymous):

is that right?

OpenStudy (turingtest):

yes it is :) sorry for the late reply